# Agent-based modeling for realistic reproduction of human mobility and contact behavior to evaluate test and isolation strategies in epidemic infectious disease spread

### v1

@@ 2.8 Stochasticity and random number generator
In order to realize stochastic events in the MEmilio ABM, we use a pseudo random number generator (PRNG). From a technical side, the choice of PRNG is important and the PRNG implementation has to satisfy several constraints. The PRNG has to be of high quality, lightweight, and lead to reproducible results for sequential and parallelized runs. A single global PRNG is inefficient in a multithreaded program because of the synchronization required. Thus, many parallel PRNGs are required that are efficiently initialized and that together generate a high quality sequence of random samples. If every agent has its own PRNG, the result of a parallelized simulation is deterministic even if the order of computation changes, e.g., due to scheduling or different thread counts. Counter-based PRNGs (cbPRNG) fulfill these requirements very well as described by [61].
@@
For a counter of size $B$ bits, the random sequence of $N=2^{B}$ samples can be split into $L$ subsequences of $b$ bits with $n=2^{b}=N/L$ samples by starting the counter for the $i$-th subsequence at $i\cdot n$ for $i=0,...,L-1$. Thus, the cbPRNG is trivially parallelizable. We assign each agent one subsequence. For each agent, we store only the counter $\widetilde{c}_{i}$ in its own local subsequence. To get the corresponding counter $c_{i}$ in the global sequence from the local counter, we combine the local counter with the index of the agent as $c_{i}=\widetilde{c}_{i}+i\cdot n$, efficiently implementable using bitwise operations. The subsequence index $i$ could be determined from the storage and iteration order of the agents. But in this model, we already store this index for other reasons. No additional storage is required either way. The key is stored only once. Only $b$ bits of storage per agent are required as opposed to regular PRNGs where each agent would need to store the full state, which is at least the size of the samples generated, more for high quality generators.
We are using Threefry-2$\times$32 implemented in the Random123 library [62] and described in [61]. This cbPRNG uses 64 bit keys and counters and efficiently generates 64 bit samples of sufficient statistical quality. Currently, we split the sequence into $2^{32}$ subsequences (i.e., up to roughly 4.3 billion agents) of $2^{32}$ samples. These numbers can be tuned in the future to support more agents, optimize storage, generate longer sequences, or increase the statistical quality further.
@@ 2.9 Parallelization
We optimized our ABM simulation using parallelization in two different ways by employing both shared and distributed memory approaches. First, in order to optimize a single simulation run we utilized multicore shared-memory parallelism on single node machines, e.g., consumer laptops. Second, we parallelized simultaneous independent runs, e.g., for calibration and uncertainty analysis on several cluster nodes using multicore as well as multinode parallelism combining both OpenMP and MPI.
@@
For multiple independent concurrent simulations, we run these in parallel either on a single node using OpenMP or over several nodes using MPI in combination with OpenMP. As we use nested parallelism within OpenMP, we specify first how many simulations to run in parallel on one node. The number of parallel simulations per node may be bounded by the available memory. The remaining cores can be used to parallelize single simulations, as described above. Finally, when using multiple nodes, we distribute the simulation runs across nodes using MPI. On each node, simulations are run in parallel using OpenMP as described above. Initial setup and final data gathering requires communication between nodes. Simulations themselves are independent of each other and do not require any communication. Thus we are able to complete several thousand runs for the calibration and for uncertainty analysis in a matter of hours, e.g., we ran 85 5368553685\,53685 536 simulation runs parallelized on 27 nodes (3 456 cores) of JURECA-DC [63] in less than eight hours.
@@ 3.1 Performance and scaling
Figure 7: Node-level weak scaling of the MEmilio ABM with different numbers of agents per processor. Runtime per time step.
Table 3: Node-level weak scaling, 200k agents per core. Runtime per time step.
@@
In Fig. 7, we provide node-level weak scaling results for our ABM with different numbers of agents per processor, with detailed timings in Table 3. We see that the scaling significantly drops from 8 to 32 cores, however, still having 33 % efficiency on the 32-fold of the initial problem. Future research will consider memory saturation issues, optimizing load balancing, and improving synchronization.
@@ 3.2.1 Model setup and parameter overview
Variable Description Value(s) or Distribution Source and explanation $v_{\textrm{P}}$ Viral load peak $8.1$ Motivated by  [44] $v_{\textrm{I}}$ Viral load incline $2$ Motivated by  [44] $v_{\textrm{D}}$ Viral load decline $-0.17$ Motivated by  [44] $\alpha$ Viral shed parameter $-7$ Motivated by  [44] $\beta$ Viral shed parameter $1$ Motivated by  [44] $s_{\textrm{f}}$ Viral shed factor Gamma($1.6,1/22$) Motivated by  [52, Fig. 3] $\lambda$ Infection rate from viral shed $1.596$ Grid search $p_{\textrm{Sym}}$ Chance to develop symptoms from an infection $\{0.5,0.55,0.6,0.7,0.83,0.9\}$ [20, Tab. 2] $p_{\textrm{Sev}}$ Chance to develop severe symptoms from a symptomatic infection $\{0.02,0.03,0.04,0.07,0.17,0.24\}$ [64] $p_{\textrm{C}}$ Chance to develop critical symptoms from a severe infection $\{0.1,0.11,0.12,0.14,0.33,0.62\}$ [65, Tab. 2] $p_{\textrm{D}}$ Chance to die from a critical infection $\{0.12,0.13,0.15,0.26,0.4,0.48\}$ [65, Tab. 2] $t_{\textrm{E}}^{\textrm{N}}$ Time from Exposed to nonsymptomatic LogNormal($4.5,1.5$) [20, Tab. 1] and references within $t_{\textrm{N}}^{\textrm{Sym}}$ Time to develop symptoms after infection in case of a symptomatic infection LogNormal($1.1,0.9$) [20, Tab. 1] and references within $t_{\textrm{N}}^{\textrm{R}}$ Time to recover in case of an asymptomatic infection LogNormal($8.0,2.0$) [20, Tab. 1] and references within $t_{\textrm{Sym}}^{\textrm{Sev}}$ Time to develop severe symptoms in case of a severe infection LogNormal($6.6,4.9$) [20, Tab. 1] and references within $t_{\textrm{Sym}}^{\textrm{R}}$ Time to recover in case of a symptomatic infection LogNormal($8.0,2.0$) [20, Tab. 1] and references within $t_{\textrm{Sev}}^{\textrm{C}}$ Time to develop critical symptoms in case of a critical infection LogNormal($1.5,2.0$) [20, Tab. 1] and references within $t_{\textrm{Sev}}^{\textrm{R}}$ Time to recover in case of a severe infection LogNormal($18.1,6.3$) [20, Tab. 1] and references within $t_{\textrm{C}}^{\textrm{D}}$ Time to die in case of death LogNormal($10.7,4.8$) [20, Tab. 1] and references within $t_{\textrm{C}}^{\textrm{R}}$ Time to recover in case of a critical infection LogNormal($18.1,6.3$) [20, Tab. 1] and references within $m_{\textrm{r}}$ Mask protection factor for the receiving person $0.25$ Motivated by  [5, 66] and manual adaptation $m_{\textrm{t}}$ Mask protection factor for the transmitting person $0.25$ Motivated by  [5, 66] and manual adaptation $m_{\textrm{c}}$ Overall mask compliance per location type $\{0,-0.1,-0.1,-0.3,-0.2\}$ Manual adaptation $p_{\textrm{f}}$ Protection from vaccination against severe infection $0.8$ Motivated by  [6] and manual adaptation $q_{\textrm{e}}$ Quarantine efficiency $0.5$ Manual assumption $q_{\textrm{d}}$ Quarantine length $10$ Manual assumption Test Sensitivity $0.71$ [67] Test Specificity $0.996$ [67] $p_{\textrm{s}}$ Testing probability for symptomatic agents $0.02472$ Grid search $\mu_{\textrm{ns}}$ Testing ratio symptomatic to nonsymptomatic agents $4.83$ Grid search Relative increased testing ratio during lockdown $0.2$ Motivated by [68] and manual adaption $d$ Initial dark figure $4.171$ Grid search $r_{\textrm{L}}$ Relative contact reduction during lockdown $0.2725$ Grid search Reduced trips during lockdown (Basic Shop, Social Event) $0.5$ Motivated by  [5, 66] and manual adaptation Reduced trips factor during lockdown (Work) $0.7$ Motivated by  [5, 66] and manual adaptation Reduced trips factor during lockdown (School) $0.0$ Schools closed Relative contact reduction after lockdown $0.5$ Manual adaptation due to reduction in positive tests. Reduced trips factor after lockdown (Basic Shop, Social Event) $0.8$ Motivated by  [5, 66] and manual adaptation Reduced trips factor after lockdown (Work) $0.75$ Motivated by  [5, 66] and manual adaptation Reduced trips factor after lockdown (School) $0.5$ Motivated by alternating lessons $\psi$ Seasonality factor April $0.95$ Motivated by  [5, 66] and manual adaptation $\psi$ Seasonality factor May $0.85$ Motivated by  [5, 66] and manual adaptation $r_{\textrm{E}}$ Percentage of the population celebrating Easter $0.2$ Motivated by  [69] and manual adaptation
@@
Table 4: Final ABM parameters used after the fitting process. For the first four parameters, values are given for the different age groups in ascending order. The values for mask compliance are given for Home, School, Work, Social Event, and Basic Shop.
Figure 8: Simulated outcomes of the fitted ABM for Brunswick between March 1st and May 30, 2021. Median simulation results are provided with a solid blue line, percentiles p25 and p75 as well as p5 and p95 are shown through differently shaded, blue regions. Reported data shown in red. Cumulative deaths (top left), cumulative detected (top right), ICU occupancy (bottom left) and newly detected infections (bottom right) are shown as outcomes.
@@
Most virus specific parameters are available through studies and literature: All viral load parameters are taken from [44], except for the distribution of individual viral sheds, $s_{f}$, which is estimated from [52, Fig. 3].
We use the LogNormal-distributed infection state times in infection states $t_{\textrm{E}}^{\textrm{N}}$, $t_{\textrm{N}}^{\textrm{Sym}}$, $t_{\textrm{N}}^{\textrm{R}}$, $t_{\textrm{Sym}}^{\textrm{Sev}}$, $t_{\textrm{Sym}}^{\textrm{R}}$, $t_{\textrm{Sev}}^{\textrm{C}}$, $t_{\textrm{Sev}}^{\textrm{R}}$, $t_{\textrm{C}}^{\textrm{D}}$ and $t_{\textrm{C}}^{\textrm{R}}$ that were also used in [20, Tab. 1]. For the probability to develop symptoms, $p_{\textrm{Sym}}$, we similarly refer to [20, Tab. 2].
@@
The probability to die, $p_{D}$, is determined with the help of the multi-center study [65, Tab. 2], where a distinction is being made between death rates of non-ICU patients and ICU patients. However, in our model, critically infected agents will always receive ICU treatment, and only these agents may die. This is not reflecting agents dying before receiving ICU treatment or with the lack of available ICU spots, which would still be listed as severely infected agents. To account for this fact, we use the death rate for ICU patients only in our model and adjust the probability to develop critical symptoms, $p_{C}$, accordingly, such that the total death rate is recovered. In the shown graphs, this artificial increase in critically infected agents is reversed, so that the correct amount of both severely infected agents and critically infected agents is recovered.
@@ 3.2.2 Model initialization
We use the data set of [38] with 373 378 agents, thus also including incommuters additionally to the base population living in Brunswick. We use the officially reported case data from [70] as well as reported ICU admissions [72]. As the officially reported case data does not report on an individual basis, but aggregates on local geographic and age group levels, and as death numbers are added only retrospectively to the case numbers, exact death days are not known. Therefore, we extrapolate the day of death from the day of infection and our model parameters. We assume the reporting date $t_{\textrm{rep}}$ of an infection to be close after symptom onset and thus estimate time of death as
@@
Within the three months period, the city of Brunswick went into a lockdown when its incidence went above 100 [70]. Before and after the lockdown, several lighter interventions like alternating lessons or remote work recommendations were in place. With the lockdown, stricter interventions such as school closures and contact restrictions had been implemented. These measures have been adapted to our model and are used in the lockdown scenarios. The lockdown spans from March 30th to April 29th, 2021 [68, 74]. Over the whole time, we have a maximum of 50 % of pupils going to school, while in the lockdown period schools are completely closed. We model a reduction of 20 % of persons going to shops and social events, increased to 50 % reduction during the lockdown. We further assume a 25 % reduction in work-related trips due to remote work, increased to 30 % during the lockdown.
@@ 3.2.3 Parameter fitting
In order to fit the parameters to the data, we use an exhaustive grid search to minimize a weighted mean squared error (MSE). In this weighted MSE, we compare the reported versus simulated data for the cumulative number of COVID-19 related deaths, the persons treated in ICU on a daily basis as well as the cumulative positive tests over the whole 90 days of the simulation. The sum of each MSE is added by an 10000:1000:3 weight ratio, respectively, addressing the different magnitudes of the time series. This ratio leads to a reasonable fit to all three data sets and was chosen through an iterative process. As the complete search space for even a small number of parameters in ABMs is often tremendously high, we focused on key parameters that could not be estimated well from the literature. We have identified five key parameters to fit, which are the linear infection coefficient $\lambda$, the initial dark figure $d$ which describes the ratio of all cases to the detected cases at the beginning of the simulation, the contact reduction when the lockdown happens $r_{\textrm{L}}$, the testing rate of symptomatic persons $p_{\textrm{s}}$ and the ratio of asymptomatic testing probability compared to symptomatic testing probability $\mu_{\textrm{ns}}$; see also Table 4. eneral, we run every parameter set 11 times and a to address the stochastic nature of our model verage the MSE to as. For simplicity, we chose an odd number to have a direct representation of the median. From these outcomes, the best average MSE is selected as the best fit.
@@
After finalization of the fitting process, we conduct 128 simulations of the fitted ABM and provide the results on cumulative deaths and detected infections as well as ICU occupancy and daily detected new infections in Fig. 8.

### v2

@@ 2.8 Stochasticity and random number generator
Stochasticity plays a major role in accurately modeling epidemic dynamics, as infection transmission and disease progression inherently involve random processes. Therefore, it must be carefully integrated to ensure realistic simulations. In order to model stochastic events in the MEmilio ABM, we use a pseudo random number generator (PRNG). From a technical side, the choice of PRNG is important and the PRNG implementation has to satisfy several constraints. The PRNG has to be of high quality, lightweight, and lead to reproducible results for sequential and parallelized runs. A single global PRNG is inefficient in a multithreaded program because of the synchronization required. Thus, many parallel PRNGs are required that are efficiently initialized and that together generate a high-quality sequence of random samples. If every agent has its own PRNG, the result of a parallelized simulation is deterministic even if the order of computation changes, e.g., due to scheduling or different thread counts. Counter-based PRNGs (cbPRNG) fulfill these requirements very well as described by [71].
@@
For a counter of size $B$ bits, the random sequence of $N=2^{B}$ samples can be split into $L$ subsequences of $b$ bits with $n=2^{b}=N/L$ samples by starting the counter for the $i$-th subsequence at $i\cdot n$ for $i=0,...,L-1$. Thus, the cbPRNG is trivially parallelizable. We assign each agent one subsequence. For each agent, we store only the counter $\widetilde{c}_{i}$ in its own local subsequence. To get the corresponding counter $c_{i}$ in the global sequence from the local counter, we combine the local counter with the index of the agent as $c_{i}=\widetilde{c}_{i}+i\cdot n$, efficiently implementable using bitwise operations. The subsequence index $i$ could be determined from the storage and iteration order of the agents. But in this model, we already store this index for other reasons. No additional storage is required either way. The key is stored only once. Only $b$ bits of storage per agent are required as opposed to regular PRNGs where each agent would need to store the full state, which is at least the size of the samples generated, more for high-quality generators.
We are using Threefry-2$\times$32 implemented in the Random123 library [72] and described in [71]. This cbPRNG uses 64-bit keys and counters and efficiently generates 64-bit samples of sufficient statistical quality. Currently, we split the sequence into $2^{32}$ subsequences (i.e., up to roughly 4.3 billion agents) of $2^{32}$ samples. These numbers can be tuned in the future to support more agents, optimize storage, generate longer sequences, or increase the statistical quality further.
@@ 2.9 Parallelization
As agent-based models typically scale with the number of agents, an efficient implementation is essential to ensure computational feasibility. Furthermore, we can make use of parallelization to significantly reduce the simulation runtime, enabling the analysis of larger and more complex scenarios. We optimized our ABM simulation using parallelization in two different ways by employing both shared and distributed memory approaches. First, in order to optimize a single simulation run we utilized multicore shared-memory parallelism on single node machines, e.g., consumer laptops. Second, we parallelized simultaneous independent runs, e.g., for calibration and uncertainty analysis on several cluster nodes using multicore as well as multinode parallelism combining both OpenMP and MPI.
@@
For multiple independent concurrent simulations, we run these in parallel either on a single node using OpenMP or over several nodes using MPI in combination with OpenMP. As we use nested parallelism within OpenMP, we specify first how many simulations to run in parallel on one node. The number of parallel simulations per node may be bounded by the available memory. The remaining cores can be used to parallelize single simulations, as described above. Finally, when using multiple nodes, we distribute the simulation runs across nodes using MPI. On each node, simulations are run in parallel using OpenMP as described above. Initial setup and final data gathering requires communication between nodes. Simulations themselves are independent of each other and do not require any communication. Thus, we are able to complete several thousand runs for the calibration and for uncertainty analysis in a matter of hours, e.g., we ran 85 5368553685\,53685 536 simulation runs parallelized on 27 nodes (3 456 cores) of JURECA-DC [73] in less than eight hours.
@@ 3.1 Performance and scaling
Figure 7: Node-level weak scaling of the MEmilio ABM with different numbers of agents per processor. Runtime per time step for up to 32 cores.
In Fig. 7, we provide node-level weak scaling results for our ABM with different numbers of agents per processor, with detailed timings in Table 3. While the efficiency scales well up to four cores, it declines significantly between 8 and 32 cores, however, still achieves $33~{}\%$ efficiency for a 32-fold increase in problem size. Future research will consider memory saturation issues, optimizing load balancing, and improving synchronization.
Table 3: Node-level weak scaling, 200k agents per core. Runtime per time step for up to 32 cores.
@@ 3.2.1 Model setup and parameter overview
Variable Description Value(s) or Distribution Source and explanation $v_{\textrm{P}}$ Viral load peak $8.1$ Motivated by  [55] $v_{\textrm{I}}$ Viral load incline $2$ Motivated by  [55] $v_{\textrm{D}}$ Viral load decline $-0.17$ Motivated by  [55] $\alpha$ Viral shed parameter $-7$ Motivated by  [55] $\beta$ Viral shed parameter $1$ Motivated by  [55] $s_{\textrm{f}}$ Viral shed factor Gamma($1.6,1/22$) Motivated by  [63, Fig. 3] $\lambda$ Infection rate from viral shed $1.596$ Grid search $p_{\textrm{Sym}}$ Chance to develop symptoms from an infection $\{0.5,0.55,0.6,0.7,0.83,0.9\}$ [31, Tab. 2] $p_{\textrm{Sev}}$ Chance to develop severe symptoms from a symptomatic infection $\{0.02,0.03,0.04,0.07,0.17,0.24\}$ [74] $p_{\textrm{C}}$ Chance to develop critical symptoms from a severe infection $\{0.1,0.11,0.12,0.14,0.33,0.62\}$ [75, Tab. 2] $p_{\textrm{D}}$ Chance to die from a critical infection $\{0.12,0.13,0.15,0.26,0.4,0.48\}$ [75, Tab. 2] $t_{\textrm{E}}^{\textrm{N}}$ Time from Exposed to nonsymptomatic LogNormal($4.5,1.5$) [31, Tab. 1] and references within $t_{\textrm{N}}^{\textrm{Sym}}$ Time to develop symptoms after infection in case of a symptomatic infection LogNormal($1.1,0.9$) [31, Tab. 1] and references within $t_{\textrm{N}}^{\textrm{R}}$ Time to recover in case of an asymptomatic infection LogNormal($8.0,2.0$) [31, Tab. 1] and references within $t_{\textrm{Sym}}^{\textrm{Sev}}$ Time to develop severe symptoms in case of a severe infection LogNormal($6.6,4.9$) [31, Tab. 1] and references within $t_{\textrm{Sym}}^{\textrm{R}}$ Time to recover in case of a symptomatic infection LogNormal($8.0,2.0$) [31, Tab. 1] and references within $t_{\textrm{Sev}}^{\textrm{C}}$ Time to develop critical symptoms in case of a critical infection LogNormal($1.5,2.0$) [31, Tab. 1] and references within $t_{\textrm{Sev}}^{\textrm{R}}$ Time to recover in case of a severe infection LogNormal($18.1,6.3$) [31, Tab. 1] and references within $t_{\textrm{C}}^{\textrm{D}}$ Time to die in case of death LogNormal($10.7,4.8$) [31, Tab. 1] and references within $t_{\textrm{C}}^{\textrm{R}}$ Time to recover in case of a critical infection LogNormal($18.1,6.3$) [31, Tab. 1] and references within $m_{\textrm{r}}$ Mask protection factor for the receiving person $0.25$ Motivated by  [5, 76] and manual adaptation $m_{\textrm{t}}$ Mask protection factor for the transmitting person $0.25$ Motivated by  [5, 76] and manual adaptation $m_{\textrm{c}}$ Overall mask compliance per location type $\{0,-0.1,-0.1,-0.3,-0.2\}$ Manual adaptation $p_{\textrm{f}}$ Protection from vaccination against severe infection $0.8$ Motivated by  [6] and manual adaptation $q_{\textrm{e}}$ Quarantine efficiency $0.5$ Manual assumption $q_{\textrm{d}}$ Quarantine length $10$ Manual assumption Test Sensitivity $0.71$ [77] Test Specificity $0.996$ [77] $p_{\textrm{s}}$ Testing probability for symptomatic agents $0.02472$ Grid search $\mu_{\textrm{ns}}$ Testing ratio symptomatic to nonsymptomatic agents $4.83$ Grid search Relative increased testing ratio during lockdown $0.2$ Motivated by [78] and manual adaption $d$ Initial dark figure $4.171$ Grid search $r_{\textrm{L}}$ Relative contact reduction during lockdown $0.2725$ Grid search Reduced trips during lockdown (Basic Shop, Social Event) $0.5$ Motivated by  [5, 76] and manual adaptation Reduced trips factor during lockdown (Work) $0.7$ Motivated by  [5, 76] and manual adaptation Reduced trips factor during lockdown (School) $0.0$ Schools closed Relative contact reduction after lockdown $0.5$ Manual adaptation due to reduction in positive tests. Reduced trips factor after lockdown (Basic Shop, Social Event) $0.8$ Motivated by  [5, 76] and manual adaptation Reduced trips factor after lockdown (Work) $0.75$ Motivated by  [5, 76] and manual adaptation Reduced trips factor after lockdown (School) $0.5$ Motivated by alternating lessons $\psi$ Seasonality factor April $0.95$ Motivated by  [5, 76] and manual adaptation $\psi$ Seasonality factor May $0.85$ Motivated by  [5, 76] and manual adaptation $r_{\textrm{E}}$ Percentage of the population celebrating Easter $0.2$ Motivated by  [79] and manual adaptation
@@
Table 4: Final ABM parameters used after the fitting process. For the first four parameters, values are given for the different age groups in ascending order. The values for mask compliance are given for Home, School, Work, Social Event, and Basic Shop. Parameters not available in the literature are either estimated by a grid search based on real-world data, where feasible, or manually adjusted, potentially informed by relevant literature.
@@
Most virus specific parameters are available through studies and literature: All viral load parameters are taken from [55], except for the distribution of individual viral sheds, $s_{f}$, which is estimated from [63, Fig. 3].
We use the LogNormal-distributed infection state times in infection states $t_{\textrm{E}}^{\textrm{N}}$, $t_{\textrm{N}}^{\textrm{Sym}}$, $t_{\textrm{N}}^{\textrm{R}}$, $t_{\textrm{Sym}}^{\textrm{Sev}}$, $t_{\textrm{Sym}}^{\textrm{R}}$, $t_{\textrm{Sev}}^{\textrm{C}}$, $t_{\textrm{Sev}}^{\textrm{R}}$, $t_{\textrm{C}}^{\textrm{D}}$ and $t_{\textrm{C}}^{\textrm{R}}$ that were also used in [31, Tab. 1]. For the probability to develop symptoms, $p_{\textrm{Sym}}$, we similarly refer to [31, Tab. 2].
@@
The probability to die, $p_{D}$, is determined with the help of the multi-center study [75, Tab. 2], where a distinction is being made between death rates of non-ICU patients and ICU patients. However, in our model, critically infected agents will always receive ICU treatment, and only these agents may die. This is not reflecting agents dying before receiving ICU treatment or with the lack of available ICU spots, which would still be listed as severely infected agents. To account for this fact, we use the death rate for ICU patients only in our model and adjust the probability to develop critical symptoms, $p_{C}$, accordingly, such that the total death rate is recovered. In the shown graphs, this artificial increase in critically infected agents is reversed, so that the correct amount of both severely infected agents and critically infected agents is recovered.
@@ 3.2.2 Model initialization
We use the data set of [49] with 373 378 agents. This includes the base population that lives in Brunswick plus the incommuters and their households. We use the officially reported case data from [80] as well as reported ICU admissions [82]. As the officially reported case data does not report on an individual basis, but aggregates on local geographic and age group levels, and as death numbers are added only retrospectively to the case numbers, exact death days are not known. Therefore, we extrapolate the day of death from the day of infection and our model parameters. We assume the reporting date $t_{\textrm{rep}}$ of an infection to be close after symptom onset and thus estimate time of death as
@@
Within the three month period, the city of Brunswick went into a lockdown when its incidence went above 100 [80]. Before and after the lockdown, several lighter interventions like alternating lessons or remote work recommendations were in place. With the lockdown, stricter interventions such as school closures and contact restrictions had been implemented. These measures have been adapted to our model and are used in the lockdown scenarios. The lockdown spans from March 30th to April 29th, 2021 [78, 84]. Over the whole time, we have a maximum of 50 % of pupils going to school, while in the lockdown period schools are completely closed. We model a reduction of 20 % of persons going to shops and social events, increased to 50 % reduction during the lockdown. We further assume a 25 % reduction in work-related trips due to remote work, increased to 30 % during the lockdown.
@@ 3.2.3 Parameter fitting
Figure 8: Simulated outcomes of the fitted ABM for Brunswick between March 1st and May 30, 2021. Median simulation results with the best-fit parameters are provided with a solid blue line, percentiles p25 and p75 as well as p5 and p95 are shown through differently shaded, blue regions. Reported data shown in red. Cumulative deaths (top left), cumulative detected (top right), ICU occupancy (bottom left) and newly detected infections (bottom right) are shown as outcomes. Vertical lines indicate the beginning of the Easter week and start of the lockdown (left line), the end of the Easter week (middle line) and the end of the lockdown (right line).
In order to fit the parameters to the data, we use an exhaustive grid search to minimize a weighted mean squared error (MSE). In this weighted MSE, we compare the reported versus simulated data for the cumulative number of COVID-19 related deaths, the persons treated in ICU on a daily basis as well as the cumulative positive tests over the whole 90 days of the simulation. The sum of each MSE is added by an 10000:1000:3 weight ratio, respectively, addressing the different magnitudes of the time series. This ratio leads to a reasonable fit to all three data sets and was chosen through an iterative process. As the complete search space for even a small number of parameters in ABMs is often tremendously high, we focused on key parameters that could not be estimated well from the literature. We have identified five key parameters to fit, which are the linear infection coefficient $\lambda$, the initial dark figure $d$ which describes the ratio of all cases to the detected cases at the beginning of the simulation, the contact reduction when the lockdown happens $r_{\textrm{L}}$, the testing rate of symptomatic persons $p_{\textrm{s}}$ and the ratio of asymptomatic testing probability compared to symptomatic testing probability $\mu_{\textrm{ns}}$; see also Table 4.
In general, we run every parameter set 11 times to address the stochastic nature of our model and average the MSEs. For simplicity, we chose an odd number to have a direct representation of the median. From these outcomes, the best average MSE is selected as the best fit.
@@
After finalization of the fitting process, we conduct 128 simulations of the ABM and provide the results and parameters, where the lowest MSE was attained, on cumulative deaths and detected infections as well as ICU occupancy and daily detected new infections in Fig. 8. Here we infer the new detected infections from the reported cumulative detected infections.